OpenStudy (anonymous):

find f(g(x)) and g(f(x)) and determine whether they are inverse of each other f(x)=3x+8 and g(x)=x-8/3

OpenStudy (anonymous):

do you know how to find the inverse of a function?

OpenStudy (anonymous):

i just know you switch the variables thats about it y=3x+8 x=3y+8 im lost after that

OpenStudy (anonymous):

f(g(x))=3(x-8/3)+8=3x-8+8=3x g(f(x))=3x+8-8/3=3x-16/3

OpenStudy (***[isuru]***):

substitute g(x) instead of x in f(x) and u will get f[g(x)] = 3x substitute f(x) instead of x in g(x) and u will get g[f(x)] = 3x - 16/3 hence f[g(x)] doesn't equal to 1/g[f(x)] they are not the inverse of each othr

OpenStudy (anonymous):

yes they are not inverse

OpenStudy (anonymous):

how do u get -16

OpenStudy (anonymous):

it's -16/3 and its the result of 8-8/3

OpenStudy (***[isuru]***):

it's not -16 it's -16/3 and it's comes this way g[f(x)] = f(x)-8/3 hence f(x) = 3x+8 g[f(x)] = 3x +8 -8/3 so -16/3 is the answer for 8 - 8/3

OpenStudy (anonymous):